OpenStudy (anonymous):

There are 8 action, 3 comedy, and 5 children’s DVDs on a shelf. Suppose two DVDs are selected at random from the shelf. Find each probability.

OpenStudy (anonymous):

1. P(2 action DVDs), if replacement occurs 2. P(2 action DVDs), if no replacement occurs. 3. P(a comedy DVD, then a children’s DVD), if no replacement occurs

OpenStudy (anonymous):

1. 8/16 * 8/16 = 1/2 * 1/2 = 1/4 2. 1/2 * 7/15 = 7/30 3. 3/16 * 5/15 = 3/48 = 1/16

OpenStudy (anonymous):

good job,ty

OpenStudy (anonymous):

you are welcome